OpenStudy (anonymous):

help completely lost... Which of the following is equivalent to log50 7 rounded to three decimal places? 2.010 0.497 –0.615 –0.854

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

Log (50^k) = log (7) K * log (50) = log (7) K = log(7) / log(50) Log 50 (7) = log 7 / log 50.

OpenStudy (dumbcow):

\[\log_{50} 7=x \leftarrow \rightarrow 7 = 50^{x}\] this tells you answer is between 0 and 1 to use calculator, apply change of base \[\log_{50}7 = \frac{\ln 7}{\ln 50} \]

OpenStudy (anonymous):

i dont have a caculator that can do that

OpenStudy (dumbcow):

sure you do...every scientific calculator has natural log (even the windows calculator :)

OpenStudy (anonymous):

oh ok thanks i willtry finding it :)

OpenStudy (anonymous):

i got 1.56228250045 for my answer and it doesnt match any of my choices

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

thanks so much i understand now

OpenStudy (dumbcow):

yw